OpenStudy (anonymous):

solve f+4=6 for f ------ g

OpenStudy (isaiah.feynman):

What grade are you in?

OpenStudy (anonymous):

f+4=6 f=6-4 f=2

OpenStudy (anonymous):

what is the dotted line over g mean?